OpenStudy (anonymous):

find the inverse f(x)= 2x+1/x-1

OpenStudy (shubhamsrg):

solve for x in terms of f(x).. that'll be the inverse.. interchange x and f(x)

OpenStudy (anonymous):

ok so x=2y+1/y-1?

OpenStudy (anonymous):

how are you by the way, long time..you are never on fb lol @shubhamsrg

OpenStudy (anonymous):

\[y= \frac{ 2x+1 }{ x-1 }\] change x for y then make y the subject. \[x= \frac{ 2y+1 }{ y-1 }\] \[(y-1)x=2y+1\] \[xy-x=2y+1\] \[xy-2y=1+x\] \[y(x-2)=1+x\]\[y=\frac{ 1+x }{ x-2 }\] so that is the answer